OpenStudy (anonymous):

if x doesn't = 2, then x^2-4/2x-4, in simplest form is? Detailed answer please

jhonyy9 (jhonyy9):

for x^2 -4 us to factoriz it formul of a^2 -b^2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

so and on denominator what can you factorizing out ?

jhonyy9 (jhonyy9):

come this is easy

OpenStudy (anonymous):

I don't know. I'm extremely confused.

jhonyy9 (jhonyy9):

why ? x^2 -4= (a-b)(a+b) when a=x and b=2 try rewriting it

OpenStudy (anonymous):

x^2-4?

jhonyy9 (jhonyy9):

yes there this have wrote you there on numerator

jhonyy9 (jhonyy9):

so you dont like cooperating anything ?

OpenStudy (anonymous):

what? I cooperate.

jhonyy9 (jhonyy9):

so than rewrite x^2 -4 factorized how will be ?

OpenStudy (anonymous):

(x-2)(x+2)

jhonyy9 (jhonyy9):

exactly

jhonyy9 (jhonyy9):

so in denominator there is 2x-4 what can you factorizing out ?

OpenStudy (anonymous):

2(x-2)

jhonyy9 (jhonyy9):

yes than now try rewriting your exercise using this facvtorized form of them ok ?

OpenStudy (anonymous):

so I would eliminate (x-2) because it is in both and I would be left with (x+2) as the answer?

jhonyy9 (jhonyy9):

so i dont know ,try rewriting these calculi here ,please

OpenStudy (anonymous):

(x-2)(x+2)/2(x-2)

jhonyy9 (jhonyy9):

yes and what is the simplified form of ?

OpenStudy (anonymous):

(x+2)/2

jhonyy9 (jhonyy9):

nice exactly hope was easy sure ,yes ?

OpenStudy (anonymous):

yes, it was easy. Thank you.

jhonyy9 (jhonyy9):

yw my pleasure good luck bye